OpenStudy (anonymous):

Consider a cost-benefit model is y=2694x/100-x, where y is the the cost for removing x percent of a given pollutant. What percent can be removed for $33,000? The answer given to me as 92.5% but dont know how to sove it

OpenStudy (anonymous):

Substitute 33,000 in for y, can you proceed from here: \[33000 = 2694x / 100 - x\]

OpenStudy (anonymous):

When I plug in 92.5% for x, which was the answer given to me by my professor it doesn't give me $33,000 it gives me 33,226. The 100-x was throwing me off too. not sure where to start

OpenStudy (anonymous):

That's due to the round off error.

OpenStudy (anonymous):

The precise answer is 92.45251303 percent. It's probably more precise than that, but when you plug that into a formula with large numbers (Such as 3 million), the error propagation gets large....

OpenStudy (gw2011):

Maybe this might help. The original equation shown by qpHalcynOn as: 2694x/(100-x)=33,000 is correct Now multiply both sides of the equation by (100-x) in order to get rid of the fraction and you get: 2694x=33,000(100-x) 2694x=3,300,000-33,000x 2694x+33,000x=3,300,000 35,694x=3,300,000 x=92.45%= approximately 92.5% (This is rounded)

OpenStudy (anonymous):

thank you :)

OpenStudy (gw2011):

You're welcome.